# QUEST#4X: an extension of QUEST#4 for benchmarking multireference wavefunction methods

### v1

@@ 4.4 Vertical Excitation Energies
[] Statistical analysis of the errors (in eV) of X-TDA, SS-NEVPT2, MS-NEVPT, SDSPT2, SDSCI, and ic-MRCISD relative to iCIPT2 for the vertical excitation energies of 110 doublet and 39 quartet states∗ Type X-TDA SS-NEVPT2 MS-NEVPT2 SDSPT2 SDSCI ic-MRCISD Doublets MD 0.17 0.07 0.08 0.08 0.02 0.02 (110 states) MAD 0.40 0.10 0.09 0.10 0.06 0.04 SD 0.58 0.15 0.12 0.12 0.08 0.06 MAX 2.42 0.63 0.41 0.44 0.44 0.32 Quartets MD -0.03 0.09 0.11 0.11 0.01 0.04 (39 states) MAD 0.31 0.11 0.12 0.12 0.05 0.06 SD 0.41 0.15 0.16 0.16 0.08 0.08 MAX 1.09 0.46 0.49 0.51 0.23 0.22 Overall MD 0.12 0.07 0.09 0.09 0.02 0.03 (149 states) MAD 0.38 0.10 0.10 0.10 0.05 0.04 SD 0.54 0.15 0.13 0.13 0.08 0.07 MAX 2.42 0.63 0.49 0.51 0.44 0.32
@@
[] Internal consistency (percentage number of states within a deviation range in absolute value) across SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, ic-MRCISD, and iCIPT2 Deviation Range (eV) SS-NEVPT2 vs MS-NEVPT2 SS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSCI SDSPT2 vs SDSCI SDSCI vs ic-MRCISD ic-MRCISD vs iCIPT2 0.05 84 % 89 % 99 % 50 % 44 % 75 % 74% 0.10 95 % 95 % 99 % 72 % 68 % 95 % 89% 0.15 97 % 96 % 100 % 82 % 81 % 100 % 95%
@@ 5 Conclusions
A comprehensive dataset, QUEST#4X, has been introduced for the purpose of calibrating multireference (MR) methods. It encompasses 110 doublet and 39 quartet excited states for the 24 radicals in the dataset QUEST#4 (which has only 51 doublet states). The iCIPT2/AVTZ vertical excitation energies reported here serve as benchmark for any other MR methods. It is of interest to note that the MR methods considered here, i.e., NEVPT2, SDSPT2, SDSICI, and ic-MRCISD, share the same function space and differ only in the determination of the expansion coefficients. Because of this, SDSCI can be taken as the initial step of ic-MRCISD and can even yield SDSPT2 and MS-NEVPT2 results automatically. Moreover, SDSPT2 also yields MS-NEVPT2 results automatically. More interestingly, the degree of agreement between SDSPT2 and MS-NEVPT2, the magnitude of deviations of SDSPT2/MS-NEVPT2 from SDSCI, as well as the iterative increments going from SDSCI to ic-MRCISD are clear indicators for the internal consistency between the methods and hence their accuracy. A good internal consistency can only be achieved by choosing appropriate active spaces for all target states, as done here. Notwithstanding this, more challenging datasets for the excited states of open-shell transition metals and -elements remain to be established. Work along this direction is being carried out at our laboratory.

### v2

@@ 4.4 Vertical Excitation Energies
[] Statistical analysis of the errors (in eV) of X-TDA, SS-NEVPT2, MS-NEVPT, SDSPT2, SDSCI, and ic-MRCISD relative to iCIPT2 for the vertical excitation energies of 110 doublet and 39 quartet states∗ Type X-TDA SS-NEVPT2 MS-NEVPT2 SDSPT2 SDSCI ic-MRCISD Doublets MD 0.17 0.07 0.08 0.08 0.02 0.02 (110 states) MAD 0.40 0.10 0.09 0.10 0.06 0.04 SD 0.58 0.15 0.12 0.12 0.08 0.06 MAX 2.42 0.63 0.41 0.44 0.42 0.30 Quartets MD -0.03 0.09 0.11 0.11 0.01 0.04 (39 states) MAD 0.31 0.11 0.12 0.12 0.05 0.06 SD 0.41 0.15 0.16 0.16 0.08 0.08 MAX 1.09 0.46 0.49 0.51 0.23 0.22 Overall MD 0.12 0.07 0.09 0.09 0.02 0.03 (149 states) MAD 0.38 0.10 0.10 0.10 0.05 0.04 SD 0.54 0.15 0.13 0.13 0.08 0.07 MAX 2.42 0.63 0.49 0.51 0.42 0.30
@@
[] Internal consistency (percentage number of states within a deviation range in absolute value) across SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, ic-MRCISD, and iCIPT2 Deviation Range (eV) SS-NEVPT2 vs MS-NEVPT2 SS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSCI SDSPT2 vs SDSCI SDSCI vs ic-MRCISD ic-MRCISD vs iCIPT2 0.05 84 % 89 % 99 % 50 % 44 % 75 % 74% 0.10 95 % 95 % 99 % 72 % 68 % 95 % 89% 0.15 97 % 96 % 100 % 82 % 81 % 100 % 96%
[] CASSCF, MS-NEVPT2, and SDSPT2 wavefunctions of excited states of H2PO and CH3 H2PO Ground state: (10a’)1(3a”)2 (3a”)2 (11a’)0 (10a’)1 (4a”)1 72% 15% 18% 62% 3% 75% 87% 3% 7% 73% 84% 5% CH3 Ground state: (3a1)2(1b2)2(1b1)1 (1b1)1 (5a1)0 (3a1)2 (1b1)1 96% 0% 0% 95% 20% 76% 75% 20% 42% 54% 54% 41% (1b1)1 (2b2)0 (1b2)2 (1b1)1 96% 0% 0% 95% 20% 76% 75% 20% 48% 48% 48% 47%
[] Vertical excitation energies (in eV) of C3H5 Typea SS-NEVPT2 MS-NEVPT2 SDSPT2 SDSCI ic-MRCISD iCIPT2b 95% intervalc AVTZ AVTZ AVTZ AVTZ AVTZ AVTZ Val. 3.57 3.56 3.50 3.28 3.36 3.40 0.02 0.03 Ryd. 5.13 5.13 5.10 4.79 4.87 4.94 0.04 0.03 Ryd. 5.70 5.70 5.68 5.35 5.44 5.46 0.06 0.03 Ryd. 5.79 5.79 5.77 5.43 5.53 5.61 0.01 0.03 Ryd. 6.12 5.78 5.93 5.74 5.79 5.71 0.05 0.02 Val. 6.20 6.20 6.17 5.73 5.86 5.98 0.02 0.02 Ryd. 7.00 6.65 6.68 6.60 6.60 6.53 0.01 0.03 Ryd. 7.92 7.92 7.89 7.41 7.54 7.68 0.02 0.02 Ryd. 7.17 7.87 7.73 7.72 7.69 7.80 0.02 0.04 Typea CI2d ic-MRCISDe MS-CASPT2f ROCC3h mcf effecti iCIPT2b 95% intervalc DZP DZP ANO-L AVTZ of UCCSD AVTZ + 2s1p on C + 1s1p on C + 1s1p1d on C + mcf: 2s2p2d Val. 3.36 3.32 3.32 3.44 0.03 3.40 0.02 0.03 Ryd. 4.70 4.68 5.11 4.94 -0.01 4.94 0.04 0.03 Ryd. 5.25 5.25 5.65 5.49 -0.02 5.46 0.06 0.03 Ryd. 5.31 5.29 5.76 5.59 0.09 5.61 0.01 0.03 Ryd. 5.50 5.46 5.73 5.65 -0.09 5.71 0.05 0.02 Val. - - 5.89g 6.15 0.04 5.98 0.02 0.02 Ryd. - - 6.36 6.20 -0.49 6.53 0.01 0.03 Ryd. - - - - - 7.68 0.02 0.02 Ryd. - - 6.90 6.79 -1.44 7.80 0.02 0.04
a
Val.: valence excitation; Ryd.: Rydberg excitation.
b
means extrapolation distance .
c
Sum of the 95% confidence intervals of ground and excited states.
d
Ref. 300; second-order MRCI with (3e,7o).
e
Ref. 301; ic-MRCISD with (3e,7o).
f
Ref. 298; ANO-L: C(14s9p4d)/H(8s4p) C[4s3p2d]/H[3s2p]; (3e,7o) for the doublets, whereas (3e,3o) for .
g
Ref. 298; CASPT2 with (3e,3o).
h
Ref. 299; mcf: molecule-centered basis functions.
i
Ref. 299; E(UCCSD/AVTZ + mcf) E(UCCSD/AVTZ).
[] CASSCF, MS-NEVPT2, SDSPT2, SDSCI, and ic-MRCISD wavefunctions of excited states of C3H5 C3H5 Ground state: (6a1)2(1a2)1(4b2)2(1b1)2 (1a2)1 (2b1)0 (1b1)2 (1a2)1 (1a2)1 (3b1)0 (1a2)1 (4b1)0 Rydberg Valence Valence Rydberg Rydberg component 41% 49% 1% 0% 1% 10% 5% 51% 28% 79% 8% 11% 41% 33% 74% 29% 25% 0% 32% 32% 43% 48% 0% 0% 0% 37% 30% 11% 14% 25% 6% 10% 71% 3% 74% 2% 3% 10% 76% 86% 41% 49% 0% 0% 0% 21% 13% 34% 26% 60% 14% 17% 55% 4% 59% 11% 12% 3% 63% 66% 40% 51% 0% 0% 0% 18% 9% 39% 28% 67% 14% 15% 52% 11% 63% 17% 16% 2% 54% 56% 36% 45% 0% 0% 0% 18% 11% 32% 23% 55% 13% 15% 48% 5% 53% 11% 11% 3% 55% 58%
@@ 5 Conclusions
A comprehensive dataset, QUEST#4X, has been introduced for the purpose of calibrating multireference (MR) methods. It encompasses 110 doublet and 39 quartet excited states for the 24 radicals in the dataset QUEST#4 (which has only 51 doublet states). The iCIPT2/AVTZ vertical excitation energies reported here serve as a benchmark for any other MR methods. It is of interest to note that the MR methods considered here, i.e., NEVPT2, SDSPT2, SDSCI, and ic-MRCISD, share the same function space and differ only in the determination of the expansion coefficients. Because of this, SDSCI can be taken as the initial step of ic-MRCISD and can even yield SDSPT2 and MS-NEVPT2 results automatically. Moreover, SDSPT2 also yields MS-NEVPT2 results automatically. More interestingly, the degree of agreement between SDSPT2 and MS-NEVPT2, the magnitude of deviations of SDSPT2/MS-NEVPT2 from SDSCI, as well as the iterative increments going from SDSCI to ic-MRCISD are clear indicators for the internal consistency between the methods and hence their accuracy. A good internal consistency can only be achieved by choosing appropriate active spaces for all target states, as done here. Notwithstanding this, more challenging datasets for the excited states of open-shell transition metals and -elements remain to be established. Work along this direction is being carried out at our laboratory.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 22373057) and Mount Tai Scholar Climbing Project of Shandong Province.
Supporting Information
Information for the radicals under concern, proper choice of zeroth-order Hamiltonian for Rydberg states, valence atomic orbitals, and CASSCF energies.
Conflicts of interest
There are no conflicts to declare.